OpenStudy (anonymous):

Explain how you can show that -3 is a solution of x + 7 = 4.

OpenStudy (anonymous):

Isolate the variable. x+7=4 -7 -7 x = -3

OpenStudy (anonymous):

That's it really. Just do everything to both sides.

OpenStudy (phi):

one way is replace x with -3 in x+7=4 replace the x with -3: -3 + 7 = 4 now do the arithmetic: -3+7 is 4, so you get 4=4 this is true, so x=-3 is a solution.

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

NP